OpenStudy (anonymous):

Anton joined a golf club two years ago. He pays an annual membership fee of $895 and a greens fee of $30 each time he plays a game of golf. The function below can be used to calculate the total yearly golfing fee, f (g), in dollars. f (g) = 895 + 30g, where g represents the number of times he played golf during the year. Last year he paid $2,065 as a total golfing fee. For how many games did he pay a greens fee? 1 points Question 2 For the function f(x) = 3x+2, find x such that f(x) = 14.

OpenStudy (hoblos):

Question 1: 2,065=895 + 30g g=39 Question 2: 14 = 3x+2 3x=12 x=4